### v1

# Neural network embeddings recover value dimensions from psychometric survey items on par with human data

Abstract
This study introduces "Survey and Questionnaire Item Embeddings Differentials" (SQuID), a novel methodological approach that enables neural network embeddings to effectively recover latent dimensions from psychometric survey items. We demonstrate that embeddings derived from large language models, when processed with SQuID, can recover the structure of human values obtained from human rater judgments on the Revised Portrait Value Questionnaire (PVQ-RR). Our experimental validation compares multiple embedding models across a number of evaluation metrics. Unlike previous approaches, SQuID successfully addresses the challenge of obtaining negative correlations between dimensions without requiring domain-specific fine-tuning. Quantitative analysis reveals that our embedding-based approach explains 55% of variance in dimension-dimension similarities compared to human data. Multidimensional scaling configurations from both types of data show fair factor congruence coefficients and largely follow the underlying theory. These results demonstrate that semantic embeddings can effectively replicate psychometric structures previously established through extensive human surveys. The approach offers substantial advantages in cost, scalability and flexibility while maintaining comparable quality to traditional methods. Our findings have significant implications for psychometrics and social science research, providing a complementary methodology that could expand the scope of human behavior and experience represented in measurement tools.
Neural network embeddings recover value dimensions from psychometric survey items on par with human data
Max Pellert1, Clemens M. Lechner2, Indira Sen3, Markus Strohmaier2,3,4 1Barcelona Supercomputing Center 2GESIS – Leibniz Institute for the Social Sciences 3University of Mannheim 4Complexity Science Hub Vienna Correspondence: max.pellert@bsc.es
1 Introduction
Psychometrics, the field of study in psychology that is crucially concerned with measurement of human attributes such as skills, personality, or values, is undergoing a revolution with the integration of novel NLP technologies to complement and extend traditional questionnaire-based workflows. Proposed ideas range from the assessment of large language models via "AI Psychometrics" Pellert et al. (2024) to "Machine Psychology" Hagendorff (2023) or, more conservatively, to use LLMs as assisting tools in human psychometrics to aid comparably cheap, in-silico questionnaire item creation Götz et al. (2023) or automated coding Nilsson et al. (2025). For the closely related task of synthetic data generation Hommel et al. (2022), LLMs have been used to simulate segments of the population to match survey results Argyle et al. (2023) over time Ahnert et al. (2025) and to replicate experimental studies Aher et al. (2023); Dillion et al. (2023).
The social and behavioral sciences crucially depend on texts that have been rated by human respondents. In psychology this is the traditional approach to analyze traits that are not directly observable such as value orientations Schwartz and Cieciuch (2022), personality John et al. (2008) and attitudes towards issues like gender beliefs Kerr and Holden (1996); Hill and Willoughby (2005); Schudson and van Anders (2022), but the same holds also for fields like political science, sociology and human-computer interaction. Questionnaires, the instruments used to collect human rater judgments for measurement, are typically developed iteratively. First, researchers compile a large set of candidate items based on theories of human experience and behavior. Items are short statements or questions believed to measure an underlying phenomenon, which humans rate on verbal and/or numeric response scales, such as a five-point agree–disagree scale. Through iterative procedures of elimination and refinement, researchers ultimately arrive at a final set of questionnaire items in which each item ideally measures only one underlying dimension of the phenomenon of interest (a property called "simple structure"). In this process, descriptive statistics and correlations among items obtained from samples of human raters are of paramount interest Repke et al. (2024). Especially the patterns of inter-correlations are crucial in identifying underlying dimensions measured by the items. However, obtaining sufficiently large samples of human raters, and letting them rate a large number of candidate items, is often costly and plagued by issues such as respondent fatigue, response style, low literacy or lack of motivation. These problems can compromise data quality and thwart a proper assessment of the questionnaire’s properties. We argue that addressing the challenge of obtaining fast, cheap and ideally unbiased information on the structure of questionnaires is a promising application area for NLP research.
Although the potential is large, critical evaluation has shown various limitations of existing attempts to marry NLP technologies with social scientific applications such as psychometric measurement Löhn et al. (2024); Sühr et al. (2024). We review selected representative approaches by other researchers and discuss the problems they face. Then, we present our novel methodological advancement of "Survey and Questionnaire Item Embeddings Differentials" (SQuID) and demonstrate with the Revised Portrait Value Questionnaire (PVQ-RR) how it allow us to overcome issues that we identified.
Research objective: We want to demonstrate the power of neural embeddings in psychometric workflows with the example of the PVQ-RR. Furthermore, we want to argue for this approach to provide useful and interesting complements to human rater judgments.
Approach: We use LLMs to create semantic embeddings of the text of psychometric questionnaire items. We choose a widely used survey on value orientations that was carefully developed and extensively validated to compare our embeddings approach to traditional human rater judgments Schwartz and Cieciuch (2022).
Results and contributions: We show that, unlike some of the related works we review claim, semantic embeddings from a general-purpose model in combination with our methodological advancement can indeed recover very well the latent dimensions of a social scientific questionnaire. According to conventionally used quantitative evaluation metrics, we achieve a performance that is at least on par with methods relying on human rater judgments. Our robust findings demonstrate the feasibility and the potential of the "semantic embeddings as human rater judgments" approach and open a large domain of interesting research questions that can be tackled in powerful novel ways by adapting well-established classical psychometric analysis pipelines. While maintaining the same level of quality, our approach based on semantic embeddings is highly flexible, cheap and scales very well.
2 Related Work
Researchers have proposed various ways of integrating LLMs in social scientific workflows. Typically, these efforts rely on repeatedly gathering stochastic model output in web interfaces of commercial, closed black-box APIs or on using older, smaller open models Binz and Schulz (2023); Fischer et al. (2023); Serapio-García et al. (2025). In this section, we review the most relevant recent literature that we found to be so far in the domain of personality questionnaires.
First uses of embeddings to test the structure of psychometric inventories
A number of recent studies have proposed using embeddings of questionnaire items from psychometric inventories—or various aggregates of items. These embedding vectors can then be used to create similarity matrices consisting of pairwise similarities between all items or item aggregates. These, in turn, can be analyzed with factor-analytic or similar dimension-reduction methods in much the same way that researchers have analyzed correlation matrices resulting from human subjects’ responses to psychometric inventories. On a fundamental level that also ensures full reproducibility, Cutler and Condon (2023) successfully replicated a classical human subject study on personality traits using an embedding approach. The authors created templates: "Those close to me say I have a [MASK][MASK] and [TERM] personality." [TERM] is inserted from a list of personality descriptors containing entries such as "condescending", the constructed statement is passed through the model and the token embeddings of the two [MASK] tokens are extracted and averaged. The results of principal component analysis on that averaged embedding were reasonably similar to results of a classical study on human ratings of personality descriptors Saucier and Goldberg (1996) with 3 out of 5 factors adequately well recovered. The authors used a less systematic and more combinatorial approach of running the same experiment with a large number of different models, all versions of BERT, to find the best-performing one according to factor congruence scores with the original human rater judgments. Interestingly, a model finetuned for Natural Language Inference (NLI) tasks, the head of which was in fact discarded for this setup (as only the two masked token embeddings are extracted), performed best. Cutler and Condon did not provide a systematic investigation on why this specific model worked best. Moreover, because of the structure of the templates containing personality descriptors, their approach is largely limited to personality inventories for which this structure is adequate. While their replication study did not provide an explicit acknowledgment of the generality of the approach, it presented pioneering evidence of the power of neural embeddings in emulating human rater judgments.
Embedding items of different questionnaires to compare them
Researchers have also taken a more macro-perspective and worked with embeddings of data sets that collect a number of different personality questionnaires. Wulff and Mata (2025) finetune a BERT based model on 200,000 pairs of items and their empirical correlations to create a domain-specific embedding model. Laudably, they share their model publicly on the Hugging Face Model Hub  [^1] https://huggingface.co/dwulff/mpnet-personality and we include it in our analyses. They use their model to embed a large number of different personality items, scales and construct names. These embeddings allow them to predict empirical similarities between items reasonably well, to detect overlaps between scales and to suggest more parsimonious taxonomies. Relying on a very similar base model, Hommel and Arslan (2025) used finetuning to create an embedding model that allows them to establish inter-item correlations, scale reliability and inter-scale correlations to some degree. Unfortunately, their model is not openly shared and can therefore not be used by us or by other researchers.
First proposals of embeddings as part of the psychometric toolbox
These approaches are generalized by Feraco and Toffalini (2025) by building on the before-mentioned "mpnet-personality" model to create a pipeline module that allows researchers to run in-silico validity checks on item candidates to identify problems in a cheap and easy way before any empirical data from human raters is collected. Similar, "pseudo-factor analysis" was recently proposed as a pre-testing approach that allows evaluating the semantic structure of personality inventories (and potentially other psychological constructs) Guenole et al. (2025). Like the approach by Feraco and colleagues, pseudo-factor analyis is based on embedding items or parcels of items and computing similarity matrices between embedding vectors, which can then be factor-analysed in the same way as inter-item correlations of personality ratings from human respondents are traditionally factor analyzed. Guenole and colleagues illustrated their approach with two widely used personality inventories, reporting that pseudo-factor analyses was able to replicate empirical findings on the structure of these inventories obtained on human subject reasonably well, at least when using (the rather permissive) oblique rotation. Despite these recent advances, it remains unclear which of the many different models and strategies for creating and combining embeddings on different items or item parcels works best, let alone why.
Missing negative correlations in embeddings of psychometric inventories
In personality inventories and other psychometric inventories, it is common practice to use both forward-keyed (e.g., "I am meticulous.") and reverse-keyed items (e.g., "I am rather careless.") to measure the same personality construct. Such item pairs would be expected to correlate negatively with each other yet load on the same factor (e.g. Conscientiousness). Likewise, constructs and their subdimensions, such as the Big Five dimensions Neuroticism and Extraversion, often correlate negatively with each other. Such negative correlations are important properties of both the semantic space of psychological constructs and the empirical data resulting from human responses to personality inventories. However, Hommel and Arslan (2025, p.7) reported an important problem: "negative coefficients are rarely produced when comparing vector representations of sentences". This is evidenced by the scatterplot of empirical and their predicted inter-item similarities using embeddings which are cutting of at 0. This problem—shared by all of the abovementioned recent embeddings-based approaches—stems from the high-dimensional embeddings of item text containing much information on general abstract linguistic features of natural language use that are similar even across otherwise semantically distinctly dissimilar item texts. The inability to reproduce negative correlations between items constitutes a major impediment for reproducing patterns of correlations and factor loadings obtained in human subjects through embeddings-based approaches. Hommel and Arslan developed a workaround that involves re-annotating an NLI corpus Williams et al. (2018) by assigning negative direction to predicted similarity when the example pair is labeled as contradictory in the corpus. They then fine-tune their model on this adapted data. This workaround seems to have the effect of increasing performance for their use case because the model better learns to handle negation and contradictions between item texts. Different to that, Wulff and Mata decide to fully drop the sign of the correlation and just consider "unsigned (absolute) correlation as a measure of item relatedness" (2025, p.1 of SI). This approach can work for the authors because all questionnaires that are analyzed by them are in the personality domain and, therefore, exhibit a substantial similarity between each other to start from. The degree of similarity still provides important information. Also, Wulff and Mata pre-process items by removing the starting pronoun "I", a form of stemming which can be expected to remove one source of confounding shared abstract linguistic features and thereby reduce similarities.
Different to these prior attempts, we aim for a general methodology that can readily be applied post-hoc to embeddings of diverse contexts also different to personality derived from any model without re-annotating training corpora and finetuning. As described in Section 3.5, we develop "Survey and Questionnaire Item Embeddings Differentials" (SQuID) as a general method that fulfills these constraints and allows to obtain negative correlations.
3 Experimental Setup
3.1 Fundamental assumptions when adapting psychometric workflows to LLMs
To empirically test their theory, Schwartz et al. collect a large number human rater judgments on a pre-specified response scale, for each of the 57 item of the PVQ-RR: $\mathbf{y}_{i}=(r_{i1},r_{i2},\ldots,r_{in})$ where each $r_{ij}$ represents the rating given by human $j$ for item $i$. In principle, this data is in the same format as the item embeddings that we will extract from the models: $\mathbf{y}_{i}=(e_{i1},e_{i2},\ldots,e_{in})$ where each $e_{ij}$ represents the element $j$ of the embedding vector for item $i$. Both types of data can be used in the same way in downstream psychometric analysis pipelines (for a schematic overview see Figure 3).
3.2 The theory of values
All the works discussed in Section 2 have so far been in the domain of questionnaires for personality traits and attendant inventories (e.g., Big Five and HEXACO), which are a natural starting point because of the long research tradition of the ’lexical approach’ investigating the semantic structure of personality-descriptive adjectives and statements Guenole et al. (2025). However, for embedding-based approaches to be a truly valuable addition to the psychometric toolkit, their viability needs to be demonstrated on additional constructs beyond personality traits. In our study, we therefore aim to go beyond personality inventories and choose a widely used and well-established survey for the important domain of values. According to Schwartz’s theory of basic human values, values are trans-situational goals that serve as guiding principles in people’s lives, expressing distinct motivational concerns. The "Revised Portrait Value Questionnaire" (PVQ-RR) Schwartz et al. (2012) contains 57 items such as "It is important to him/her to form his/her views independently.", "It is important to him/her that his/her country is secure and stable.", "It is important to him/her to have a good time.", which can be aggregated to 19 fine-grained value dimensions such as "Self-direction Thought", "Security-Societal" and "Hedonism" (3 items per dimension). Human respondents rate how similar they view themselves to the person described in the item on a 6-point scale ranging from "not like me at all" to "very much like me". The survey scoring rules that specify the dimension that each item belongs to and can be aggregated to through averaging have been developed in a theory-guided way that has been cross-culturally validated and refined since the 1980ies with human rater judgments from 49 countries Schwartz and Cieciuch (2022).
3.3 Multidimensional Scaling and the Centrality of Negative Correlations
Importantly, negative correlations between items play a very important role in Schwartz’s theory of the content and structure of human values. Schwartz’s values are structured in a circumplex model based on the compatibility and conflict between underlying motivations, whereby values that motivational oppose each other are placed on opposite positions of the circle and correlate negatively with each other (e.g. Conformity vs. Self-direction), whereas closely related values that are motivational compatible with each other are positively correlated and located close to each other on the same side of the circle (e.g. Achievement and Power). The individual values can be further aggregated to four higher-order dimensions corresponding to four segments of the circle that emerge with great regularity in data from human respondents Lechner et al. (2024).
This circular structure is uncovered using a dimension-reduction technique that is different from principal component analysis and factor analysis as typically used for personality inventories: Multi-dimensional scaling (MDS) Jung and Takane (2015); Groenen and Borg (2015). Schwartz’s theory and attendant inventories were historically developed in close relationship with this technique. MDS maps high-dimensional objects such as vectors of human rater judgments or semantic embeddings to points in a lower-dimensional, typically 2D, space. The technique works in such a way that points corresponding to similar objects are located close together, while dissimilar objects are located far apart, thereby preserving distances.
The MDS configurations that are computed from human rater judgments confirm those relationships between values as postulated by the theory. We can therefore empirically validate the theoretically assumed circular placement of values using MDS with human rater judgments Schwartz and Cieciuch (2022), while, until now, it was an open question if the same is possible with an embedding approach. The paramount importance of negative correlations between items renders Schwartz’s value theory and the PVQ-RR an ideal (though challenging) case to refine embeddings-based approaches such that they reproduce negative correlations between items foundational to value theory.
3.4 Embedding Models
The speed of development makes it impossible to provide a systematic model review that stays state-of-art long enough to even account for the length of typical publication cycles. For this reason, we instead employ strategic selection to include models in our analyses that are representative for specific model classes and types. This way, we can provide some guidance on which models typically to look for in this application area also in the future.
First, we choose the currently best performing open model for semantic textual similarity (STS) tasks in "MTEB: Massive Text Embedding Benchmark" Muennighoff et al. (2023), Linq-Embed-Mistral (7.11B, prompt, Choi et al. (2024)). In addition, we choose the currently best performing closed API model in the benchmark, gemini-embedding-exp-03-07 (parameter count not disclosed, no prompt). Furthermore, we include two smaller models created with innovative training and data strategies that allow them to perform well for their size, jina-embeddings-v3 (572M, no prompt, Sturua et al. (2024)) and KaLM-embedding-multilingual-mini-instruct-v1.5 (494M, prompt, Hu et al. (2025)). Finally, we include the model mpnet-personality (109M, no prompt) that, as discussed before, was created through domain specific finetuning. Our approach does not need extensive prompt engineering, any sophisticated token probability extraction strategies or accounting for stochasticity of outputs. In a growing social scientific literature of using LLMs for empirical investigations, instead of interacting with models on the textual output level we show the advantages of working on the fundamental level of semantic embeddings. We use the exact wording from the PVQ-RR questionnaire specifications (see Subsection A.7) for those models that allow to include a prompt [^2] We run all models on CPU to embed all 57 short texts of the items in less than 20 minutes on a ThinkPad T480.. The PVQ-RR is available in gendered version using different pronouns. We account for this feature by creating an average of the female and male item embeddings in order to reduce possible gender biases.
As we go through in detail in Subsection 4.1, we find evidence that the MTEB benchmark is actually a good guide for selecting models for this task. There is reason to expect that future models that score even better on the benchmark will likely still outperform our already strong results.
3.5 Survey and Questionnaire Item Embeddings Differentials (SQuID)
Working with raw embeddings as directly retrieved from the models, we faced the same problem of missing negative correlations as the other researchers whose work we discussed in Section 2 (see the similarity matrix on the left of Figure 4). As we will use MDS for dimensionality reduction which builds on preserving distances, it was especially important for us to find a way to make the signal of negative correlations stronger.
As discussed before, the root of the problem lies in the high-dimensionality of the item embeddings that encode many common linguistic features of natural language use which artificially boost similarity between items. To create a vector that captures this unneeded information, we take an average embedding over all items (a "questionnaire embedding") and subtract it from each of the item embeddings: $\mathbf{y}_{i}-\overline{\mathbf{y}}$ for item $i$, where $\overline{\mathbf{y}}$ is an average over all 57 items of the PVQ-RR: $\overline{\mathbf{y}}=\frac{1}{57}\sum_{i=1}^{57}\mathbf{y}_{i}$. This straightforward method is inspired by the the early examples of vector arithmetic demonstrated with Word2Vec Mikolov et al. (2013) and the more recent idea of community embeddings Waller and Anderson (2021).
As can be seen in the right hand part of Figure 4, this approach is simple yet very effective in making the signal of negative correlations appear without any domain-specific finetuning or other adaption procedures. SQuID works with any model and embeddings of arbitrary dimensions, offering a high degree flexibility. We believe that a substantial part of the under-performance of embedding approaches compared to human data in other studies, for example in Wang et al. (2024), is due to methodological reasons and can be overcome with SQuID.
AC BEC BED COI COR FAC HE HUM POD POR
gemini 0.45 0.74 0.71 0.80 0.74 0.47 0.63 0.34 0.74 0.74
jina 0.60 0.77 0.76 0.73 0.77 0.51 0.69 0.22 0.58 0.83
kalm 0.61 0.72 0.70 0.76 0.79 0.34 0.62 0.23 0.64 0.81
linq 0.72 0.89 0.68 0.83 0.90 0.53 0.72 0.46 0.82 0.87
mpnet 0.37 0.74 0.52 0.68 0.75 0.18 0.62 -0.05 0.63 0.88
$\mathcal{N}(0,\,1)$ -0.06 -0.06 -0.06 -0.04 -0.06 -0.04 -0.04 -0.04 -0.04 -0.06
human 0.60 0.74 0.69 0.75 0.81 0.68 0.73 0.47 0.73 0.76
SDA SDT SEP SES ST TR UNC UNN UNT $\mu$
gemini 0.78 0.57 0.54 0.67 0.65 0.70 0.53 0.66 0.72 0.64
jina 0.78 0.66 0.49 0.81 0.73 0.76 0.49 0.79 0.73 0.67
kalm 0.79 0.68 0.62 0.68 0.71 0.82 0.63 0.85 0.71 0.67
linq 0.82 0.68 0.66 0.91 0.78 0.83 0.77 0.92 0.85 0.77
mpnet 0.55 0.36 0.50 0.58 0.72 0.81 0.49 0.86 0.63 0.57
$\mathcal{N}(0,\,1)$ -0.06 -0.04 -0.06 -0.04 -0.06 -0.06 -0.04 -0.04 -0.06 -0.05
human 0.66 0.65 0.58 0.77 0.67 0.80 0.73 0.83 0.71 0.70
Table 1: Internal consistency as measured through Cronbach’s alpha for each of the 19 value dimensions and the average over dimensions. Rowname abbreviations refer to the models covered in Subsection 3.4 (all embeddings were SQuID treated). We report a random baseline of embeddings created from a normal distribution with mean zero and unit variance (averages over 10 000 samples). Human data is retrieved from Table S2 in SI of Schwartz and Cieciuch (2022). Linq-Embed-Mistral is generally the best performing embedding model with an average value for Cronbach’s alpha that is above even the human benchmark. Values close to 0.8 or above for most value dimension point to adequate internal consistency (with notable exceptions such as HUM/humility, which is also low in the human data). For fully spelled out dimension names in the column header see Figure 4.
4 Results
The results presented here are derived from SQuID treated embeddings from the best-performing model "Linq-Embed-Mistral".
4.1 Validity Checks
4.1.1 Internal consistency
Following Schwartz and Cieciuch (2022), we employ Cronbach’s alpha to assess the internal consistency of latent dimensions. Through functional equivalence in this context (for details see Subsection A.3), we can directly compare embeddings of all our candidate models with human data [^3] We retrieve the human values from Table S2 in SI of Schwartz and Cieciuch (2022) which reports averages over 49 countries.. Cronbach’s alpha compares the amount of shared variance among the items of a dimension to the amount of overall variance. In previous research, performance baselines of random embeddings have been shown to be higher than intuitively expected in some scenarios Wieting and Kiela (2019). To assess which baseline levels we reach, we produce a vector with 4096 normally distributed values ($M_{ij}\sim\mathcal{N}(0,\,1)$) for each item. We repeat the procedure 10 000 times and report average values of Cronbach’s alpha for each dimension in these random embeddings. In addition to reporting all 19 values for each model, human data and random embeddings, we also provide an overall average over all dimensions. In Table 1, we find that Linq-Embed-Mistral is the best performing approach. It beats the other approaches 14 out of 19 times and has the highest average Cronbach’s alpha over all dimensions (including being slightly higher than the human data).
4.1.2 Dimension-dimension correlations of human rater judgments and embeddings
We create a dimension-dimension similarity matrix using Pearson correlation for both the embeddings (shown in Subsection A.6) and the human data. The 171 pairs of both types of data positively correlate substantially: $0.74$ [$0.66,0.80$]. Figure 1 shows the scatter plot and parameters of a simple linear model that explains $55\%$ of the variance observed.
Figure 1: Relationship between item similarity (correlation) derived from embeddings (SQuID treated) and human data. We explain more than half of the variance with a simple linear model.
4.2 MDS Configuration from Embeddings
We turn the dimension-dimension correlation matrix of Subsection A.6 into a dissimilarity matrix. We use multidimensional scaling444mds with ’type=”ordinal”’ from the R package smacof Mair et al. (2022), for details see code file. to reduce the high-dimensional dissimilarity to a two dimensional space that preserves distances between dimensions. Figure 5 shows the resulting circular structure.
4.3 Procrustes Matrix Transformation
Next, we want to assess quantitatively the correspondence of the embeddings MDS configuration to the empirical configuration of Schwartz and Cieciuch (2022) [^5] We compute the human MDS configuration from the pooled correlation matrix over all 49 countries shared by the authors as Table S10 of SI. derived from human rater judgments. As MDS configurations are not unique, arbitrary matrix transformations such as translation, rotation and reflection are admissible because these operations do not change pairwise distances. We use Procrustes matrix transformation666procrustes from the R package vegan with the human configuration as target and the embeddings configuration as testee. This way, we transform the embeddings configuration to achieve maximum similarity with the human rater judgments configuration.
4.4 Comparison of Configurations from Human Data and from Embeddings
To quantitatively assess the similarity of the two MDS configurations after Procrustes transformation of the embeddings MDS configuration matrix, we compute cosine similarities of the vectors of the two first dimensions (X coordinates in the configuration plot) as well as the two second dimensions (Y coordinates). These congruences coefficients are $0.88$ for the first and $0.82$ for the second dimension. To further assess similarities and differences between individual dimensions, Figure 2 allows for visual assessment of the two configurations.
Figure 2: Comparison of MDS configurations from embeddings and from human data. The configurations are shown after Procrustes similarity transformation. Human data in red, embeddings in blue. We find few long and few strongly overlapping lines evidencing general similarity of the two configurations derived from two distinct types of data. For fully spelled out dimension names see Figure 4.
5 Discussion
Our validation checks uncover that value dimensions contained in neural embeddings can display internal consistency on par with those derived from human data (or even above). As a comparison baseline, we find no meaningful values for randomly initiated embeddings. Correlation of dimension-dimension similarities for human data and embeddings show substantial agreement between the two types of data. We find no bias in the regressing one on the other, a strong positive relationship and a moderately high coefficient of determination for a simple linear model.
In the embeddings configuration of Figure 5, we can already see the circular structure emerging that is postulated by Schwartz and Cieciuch (2022). Values that are facets of the same phenomenon (like "Universalism-Concern", "Universalism-Nature" and "Universalism-Tolerance") are placed close together. Values that are seen as distant to Universalism (such as "Power Dominance" and "Power Resource") are placed far apart on the circular structure, roughly on the opposite side. The same can be observed with the facets of "Self-Direction" in comparison to the facets of "Conformity". The results allow for a visual interpretation that follows quite closely along the lines of the canonical theory of value orientations that powers the PVQ-RR.
Comparing the two MDS configuration after Procrustes transformation shows fair similarity according to conventional interpretations of congruence coefficients (for details see Section A.5). That we find the two configurations to be fairly similar on this formal criterion is quite surprising given that they derive from two very distinct data sources. Inspection of individual value dimensions in Figure 2 reveals strong overlaps of the two solution with the facets of Security (Security Personal and Security Societal) and basically identical placement of some dimensions such as Conformity-Interpersonal, Power Resources and Universalism-Tolerance. While other are placed closely like Self-direction Thought and Power Dominance, dimensions can also be distinctly different between the two configurations, for example Stimulation (ST) that is placed much closer to Hedonism (HE) in the embeddings configuration.
6 Conclusion
We introduced our methodological advancement of SQuID and show with the example of the PVQ-RR that it allows an embedding approach to reach performance on par with human data in extracting latent value dimensions according to several evaluation metrics.
We envision many new directions of research to emerge from psychometric approaches to AI. As one example, our approach allows for a systematic and rigorous investigation of potential cultural biases contained in LLMs: Schwartz and Cieciuch (2022) confirmed their theories of value orientations with data from 49 different cultures using translated questionnaires. We see an exciting opportunity to use our approach and MDS configurations on the level of those individual countries to systematically assess closeness or distance of an LLM to certain cultural value orientations quantitatively by computing congruence coefficients and other error metrics. Generally, working on the embedding level as we propose is limiting sources of potential methodological biases stemming from token extraction strategies that are necessary when working with generated text instead.
We view our proposal as complimentary to both traditional workflows as well as to novel approaches in the (social) sciences. By adopting embeddings approaches, we don’t aim to replace human participants. Also well into the future, researchers will have very good reasons to work with humans, at the bare minimum to provide the necessary reference points for comparison as we demonstrated in our analyses. But different to ideas discussed before, such as "pseudo-factor analysis" Guenole et al. (2025), we want to go beyond merely creating another step in the typical validation pipeline in psychometrics. We see potential beyond that: With these novel technologies we can potentially investigate the full and diverse spectrum of expression in all natural languages. This creates novel opportunities to expand the scope of human behavior and experience represented in measurement tools.
Limitations
As mentioned in several places in the manuscript, future research will have to uncover more of the peculiarities of that novel kind of data in the form of semantic embeddings for (social) scientific applications. More experience with the data will in turn help the methods (and the theories). We see "neural embeddings as human rater judgments" as a good working hypothesis and don’t want to imply that both data are exactly the same and capture human behavior and experience in exactly the same way. As with other synthetic data, novel treatments such as specific transformation (in addition to general methodologies like SQuID) and other processing steps may be necessary depending on the application context.
Questions of measurement validity arise: Do findings from the human domain hold generally? In this emerging field, we don’t yet have established answers to these questions. We can face unexpected problems, as for example with questionnaire scoring rules that do prescribe to reverse certain items for data quality reasons. With human rater judgments this can be done in a straightforward way (by subtracting the rating from the maximum score possible as a response), but how do we "reverse" semantic embeddings in a meaningful way? Novel methodological questions like these appear often and have to be addressed in future research. With our work, we want to make a strong argument that it is worth attempting to provide answers.
The PVQ-RR is openly available on the internet and it could have been part of the training corpus of the models that we use. The models’ direct memorization of its items and the underlying theory could potentially influence our results and limit the scope of our methodology when applied to different, less easily accessible contexts. We don’t think that this is a crucial issue as we provide no information to the models on our aim to reconstruct value configurations from embeddings, by using only the generic prompt from the survey instructions (see Subsection A.7) or by using no prompt at all. Furthermore, our best performing model "Linq-Embed-Mistral" has only moderate size with 7.11B parameters and performs well in the diverse tasks of the "Massive Text Embedding Benchmark" Muennighoff et al. (2023), which points to its general capabilities.
The notion of "similarity" is actually not well-defined as a classical discussion in philosophy shows. The argument goes against the existence of a strong logical underpinning of notions of similarity, while it’s practical relevance is acknowledged. Similarity is considered "still serviceable in the streets" Goodman (1972). Some parts of the original logical argument have been debated and countered Leitgeb (2007). We further want to note that our metaphor of semantic embeddings as human rater judgments helps us to prevent some concerns: Each dimension of semantic embeddings as the equivalent to a human rater judgment provides the necessary context ("similar to what?") to justify relative assessments of similarity for our use case.
Ethics Statement
The technology of using semantic embeddings in research in ways that are similar to human rater judgments can become a common and established practice (as proposed for example by Guenole et al., 2025). For researchers, we see obvious benefits as they will be provided with a novel way of gathering and analyzing data. This approach potentially promises to capture a richer tapestry of human behavior and experience than conventional methods relying on WEIRD (Western, Educated, Industrialized, Rich, Democratic) college student populations. In addition to such convenience samples, human rater judgments are typically sourced from crowdworkers that carry out tasks at often low hourly wages. There is reason to hope that crowdworking platforms react to technology changes with a quality initiative that leads to better opportunities for crowdworkers to earn adequate incomes.
Still, crowdworking delivers incomes to people around the world that come from many marginalized groups. Although it can leave workers underpaid, it is an integral and existential part of many lives. Failures of platforms to react and dwindling numbers of crowdworking jobs (that are instead outsourced to LLMs) can leave many people worse off than before.
On another, more distant future, level, we could also see some danger in people accepting analyses based on semantical embeddings as some kind of "gold standard", a LLM normality to compare with the results of people, forgetting about many possible biases. Abstractly, like any assessment, this could increase the danger of diagnosing somebody as abnormal vis-a-vis an unattainable standard.
Generally, we see potential harm to fall mostly proportional on all parts of the population in terms of race, gender or class. However, the lack of jobs for crowdworkers can be expected to have a disproportional affect on people on low incomes in less affluent (and predominantly English-speaking) parts of the world. In addition, we try to limit the potential harm of creating novel synthetic LLM norms by restricting ourselves to a questionnaire of value orientations such as the PVQ-RR that is rather uncontroversial and is not positing any "correct" or "normal" value orientations for individuals.
Appendix A Appendix
In our paper repository [^7] https://anonymous.4open.science/r/embeddings_values-BFBE/, we share documented code that allows other researchers to fully replicate our results using openly available data.
A.1 Older lines of research on survey item similarity
For completeness sake, we want to mention that work on semantic similarity of test items goes back before the recent developments in NLP research of 2017/18, e.g. Arnulf et al. (2014). The usual approach back then was to create item embeddings and to compare them with each other for similarity using latent semantic analysis (LSA) Rosenbusch et al. (2020); Arnulf and Larsen (2021). Schematically, researchers used singular value decomposition on the co-occurence matrix of terms within items. The creation of semantic networks of items is a similar, related approach Christensen and Kenett (2021). Classic use cases for those approaches involve researchers to come up with sample items that are then used to find already existing scales containing similar items or to check for item redundancy.
A.2 Different, but related uses of semantic embeddings as human rater judgments
Researchers have used embeddings that they trained to predict rating scales as a way to analyze descriptive wordresponses by human participants Kjell et al. (2022). They find that a BERT model works well for this task, especially with word embeddings. With a slightly different twist, researchers used item embeddings together with human rater judgments of personality items to predict ratings of completely new items that had not been rated by any participant yet Abdurahman et al. (2024).
A.3 Cronbach’s Alpha of human rater judgments and embeddings
We compute Cronbach’s Alpha according to the usual formula
$\alpha={k\over k-1}\left(1-{\sum_{i=1}^{k}\sigma^{2}_{y_{i}}\over\sigma^{2}_{y}}\right)$
where $k$ represents the number of items in the dimension, for the PVQ-RR $k=3$
$\sigma_{y_{i}}^{2}$ the variance associated with each item $i$
$\sigma_{y}^{2}$ the variance associated with the total scores $y=\sum_{i=1}^{k}y_{i}$
For the human rater judgments:
$\mathbf{y}_{i}=(r_{i1},r_{i2},\ldots,r_{in})$
where each $r_{ij}$ represents the rating given by human rater $j$ for item $i$.
For the embeddings:
$\mathbf{y}_{i}=(e_{i1},e_{i2},\ldots,e_{in})$
where each $e_{ij}$ represents the element $j$ of the embedding vector for item $i$.
A.4 PCA instead of MDS
As a robustness check, replacing MDS with PCA888prcomp in R with ’center = TRUE’ and ’scale = TRUE’ and plotting the first two dimensions shows similar clustering to the MDS solution but less circular structure as PCA does not guarantee to preserve distances:
A.5 Congruence coefficients
Figure 4: Effect of questionnaire embedding subtraction on the item similarity matrix. Left hand figure shows the raw embeddings, right hand figure the SQuID treated embeddings. By using SQuID to normalize for common standard features of natural language use, negative correlations appear.
Labels refer to Embeddings and human rater judgments from Schwartz Schwartz and Cieciuch (2022). Congruence coefficients are computed for each of the two dimensions after transformation of the embeddings configuration. There is no strict standard, but according to typical interpretations, values in the range of $0.85\text{-}0.94$ between two compared configurations display fair similarity Lorenzo-Seva and Ten Berge (2006). To establish significance, we create 2 times 1000 random vectors, use Procrustes transformation to maximize similarity between them and compute their cosine similarity. We repeat this procedure 1000 times and find $p<0.001$ for the cosine similarities we report above over all runs, indicating that they are significantly different from a random baseline and cannot be explained as an artifact of Procrustes transformation.
A.6 Effect of Questionnaire Embedding Subtraction
Figure 4, from left to right, shows that through SQuID we achieve negative correlations in the item-item similarity matrix (sorted alphabetically by dimensions). The similarity matrix of the same SQuID treated embeddings averaged by dimension:
We compute correlation coefficients between SQuID treated embeddings that are averaged by dimension. We don’t show the diagonal of 1s to better allow to observe correlation ranges.
A.7 Prompt
For all models that allow to include a prompt, we use the following wording taken exactly as-is from the specification of the English version of the PVQ-RR: "Here we briefly describe different people. Please read each description and think about how much that person is or is not like you. Put an X in the box to the right that shows how much the person described is like you."
Figure 5: Unrotated MDS configuration derived from SQuID treated embeddings. As value dimensions are placed according to their distances to each other, a circular structure emerges. For fully spelled out dimension names see Figure 4.